# Reviewing Intelligent Cinematography: AI research for camera-based video production

### v1

@@ Abstract
This paper offers a comprehensive review of artificial intelligence (AI) research in the context of real camera content acquisition for entertainment purposes and is aimed at both researchers and cinematographers. Considering the breadth of computer vision research and the lack of review papers tied to intelligent cinematography (IC), this review introduces a holistic view of the IC landscape while providing the technical insight for experts across across disciplines. We preface the main discussion with technical background on generative AI, object detection, automated camera calibration and 3-D content acquisition, and link explanatory articles to assist non-technical readers. The main discussion categorizes work by four production types: General Production, Virtual Production, Live Production and Aerial Production. Note that for Virtual Production we do not discuss research relating to virtual content acquisition, including work on automated video generation, like Stable Diffusion. Within each section, we (1) sub-classify work by the technical field of research - reflected by the subsections, and (2) evaluate the trends and challenge w.r.t to each type of production. In the final chapter, we present our concluding remarks on the greater scope of IC research and outline work that we believe has significant potential to influence the whole industry. We find that work relating to virtual production has the greatest potential to impact other mediums of production, driven by the growing interest in LED volumes/stages for in-camera virtual effects (ICVFX) and automated 3-D capture for a virtual modelling of real world scenes and actors. This is the first piece of literature to offer a structured and comprehensive examination of IC research. Consequently, we address ethical and legal concerns regarding the use of creative AI involving artists, actors and the general public, in the conclusion. Learning from the implications of Deep Fake research (mainly the non-consensual use of women and children’s faces for generating pornographic content), we emphasize the importance of awareness, responsibility and accountability regarding code sharing.
@@
Cinematography, Vision, Graphics
@@ 1 Introduction
Intelligent cinematography (IC) research leverages artificial intelligence (AI) to assist camera-based tasks in media production. For pre production, this concerns research involved with camera planning, for example using a script [213, 70] or user input [55, 272]. For production, this includes work on automated camera-control [145, 60, 158, 26, 277], methods for content acquisition [166, 196, 296, 139] and assisting directors and camera operators [73, 146, 105, 103, 36, 27]. For post production, this involves work that avoids the need for post-production, such as through in-camera virtual effects (ICVFX) [228, 125, 76, 177], or for enhancing or diminishing camera artefacts such as light-water reflections [270]. For clarity, our definition of IC encompasses research focused on real content acquisition. This excludes research aimed at assisting post production or automated content generation using diffusion models.
IC consequently influences the worlds of cinema, gaming and televised broadcasting. The symbiotic relationship between research and video production has been historically important to the progression of video entertainment. Yet, as this relationship evolves the types of content, styles of production, and thus fields of research, have branched into separate domains. For example, we find work that automates manual capture processes [70, 139] as well as work that elevates abstract goals like enhancing audience immersion [105, 257]. This variety of objectives makes it challenging to establish a unified definition of the IC research landscape, thus limits researchers and cinematographers who wish to understand the general state and trends. Moreover, there is an absence of both general review papers and more specialised articles. Hence, in this paper, we attempt to overcome these challenges by defining and presenting a structured review on the state and future of IC research.
This review paper categorizes work by production medium. Similar to the definition of an artistic medium, this groups research by the relevant tools and practices involved with the various production types, rather than grouping work by objective. For instance, live sport and event broadcasting productions share many cinematographic practices, so are grouped under the same heading, Live Production. Doing so allows us to: (1) further sub-categorize according to related AI research fields; this provides a holistic view of the IC landscape while maintaining relevance to each production medium, (2) include research topics which, while not targeted at creative video production, have direct relevance to it, and (3) offer a view of the research landscape that cinematographers can refer and relate to. The topics covered for each medium are outlined below.
@@
1.1 Scope and related work
The focus of this paper is camera-oriented research; hence we do not cover popular creative AI topics such as large language models (LLMs) for script generation or generative AI for video generation. We acknowledge, however, that these may become directly involved in cinematography in the future. For instance, LLMs may prove useful for text-driven camera control or workflow optimisation. Moreover, the rapid progression of generative AI models could deliver high quality video-based methods that offer an alternative solution to automated 3-D capture. As the relevant literature does not yet consider camera-based objectives tied to these technologies we do not investigate these paradigms in the main body. Instead, we provide background on generative AI and discuss future use cases of these technologies within the relevant subsections of the main body.
There are a few articles that review AI in the context of cinematography and those that do focus on special cases. For example, Christie et al [57] reviews camera control for computer graphics. This paradigm involves view point computation and motion planning systems, and considers technical constraints such as direct and assisted control, and cinematic constraints such as view composition and pose. Similarly, Chen and Carr [46] reviews autonomous methods for camera planning, control and selection for multi-camera systems. Furthermore, Qi et al [198] reviews camera calibration techniques. Other works look at use-cases for IC such as Galvane et al [80] for video editing, Young et al [281] and Sharma et al [229] for gaming and animation and Mademlis et al [156] for unmanned aerial vehicle (UAV) control. Our survey aims to bridge related works to provide a holistic view of the IC landscape so that computing and cinematographic communities mutually benefit from the discussion.
The structure of this paper is as follows. In Section 2 we introduce the technical background on AI research commonly found in IC. The main body is contained in Section 3, where the order of the subsections is reflected in the prior list. For each subsection, we present an overview of the current paradigms and approaches and conclude with a discussion on the current challenges and future works. Lastly, in Section 4, we offer our concluding remarks and outline research that we believe has the most potential moving forward. As cinematography typically involves capturing real people, we append a section on social responsibility to our conclusion to highlight the importance of ethically driven research and content capture. We continue to emphasize the importance of awareness, responsibility and accountability regarding code sharing throughout the paper.
@@ 2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) are a class of neural network that extract the features of an image-type by filtering through a neighbourhood of pixels with learnable filter coefficients, to learn common patterns within a set of training image(s). Regarding IC, CNNs provide backbone architectures for visual generative AI (Subsection 2.2) and object detection (Subsection 2.3).
A common application of CNNs is in image classification, where the basic structure is shown in Figure 1. First, an image is input to the first layer through $N$ channels, where each channel represents a different representation of an image. For example, an RGB image will have three values (red, green and blue) for each pixel, thus $N=3$. The outputs of this layer then feed into the hidden layers. The outputs of each layer are called feature maps. The last hidden layer transforms the feature maps to feature vectors, so that the output layer can interpret a class probability using the feature vector. An image is then classified if it’s probability falls within a predefined confidence-interval. In the case of Figure 1, if our confidence interval is $0.5$, we classify the image as containing dogs.
@@
    A convolution layer applies an $n\times m$ kernel (filter) over each pixel neighbourhood from an input channel. To deal with pixels on the perimeter of an image, we generally apply padding (the type of padding can vary). Dilated [284] or deformable convolutions [64] are also used to modify the structure of the kernel for special cases. In addition, a stride can be applied to skip some pixels, i.e. the filter applies to the pixels located $s$ distances from each-other ($s=1$ means all pixels are filtered and $s=2$ means every second pixel in a row and column is taken).
@@
    An activation function determines how the features of a layer are transformed into an output. To be used in a hidden layer, the activation function needs to be differentiable and nonlinear, otherwise we would not be able to calculate the gradients of parameters during a backward pass of the network. Rectified linear unit (ReLU) and LeakyReLUs layers are frequently used as they are a differentiable non-linear function, with the added benefit that they are not susceptible to vanishing gradients (where gradients of parameters become too small to make regressive change to the model’s parameters). Several types of activation function exist, as discussed in Baheti [12].
@@
    A fully-connected layer is one where each node is dependant on all the outputs from the previous layer. This allows the network to introduce a wide range of dependencies between parameters.
2.2 Generative AI
Generative AI concerns various domains, for instance text-to-video synthesis [14, 1, 143] and novel view synthesis for dynamic scenes [141, 196, 78]. As the novel view synthesis problem-landscape is broad and specially relevant, we reserve the discussion for a later subsection. In this subsection, we focus on generative AI involving 2-D images and video.
The purpose of generative AI varies drastically depending on its application. For example, it can involve learning representations for a specific medium, e.g. text or images, and translating/transforming the representation into alternative representations to produce related content e.g. images generated by prompts and vice-versa. Terms such as text-to-image or image-to-text generation, fall under the broader class referred to as sequence-to-sequence translation [254, 276, 183, 75]. Figure 2 demonstrates how abstractions such as text, images and video frames could be sequenced, and simplifies an encoder-decoder pipeline that uses pre-trained networks to inform the encoding, decoding and/or transformation process. Overall, there are four prominent architectures that contribute separately towards image and video synthesis: (1) encoder-decoder architectures, (2) generative adversarial networks (GANs) [63, 87, 123], (3) transformers [253] and (4) diffusion models [102, 291, 72]; discussed below.
Figure 2: Text, image and video frame input sequences are encoded into a feature/latent representation and translated/transformed into alternative representations that are decoded to output other tangible representations such as text, images and video frames. $s_{i}\in S_{j}$ illustrates how a sequence may be constructed from a set of input data. $\xi_{n}$ encodes the input sequence into a representation, $n$, and a function, $F$, translates the representation into $m$. $\xi_{m}^{-}1$ decodes $m$ back into a sequence. $\phi_{k}$ is a pre-trained network that can be used during the encoding, decoding and/or transformation stage.
In a GAN (see Brownlee [30]) a random vector is fed into a generator to produce a synthetic representation. Subsequently a comparison is made between the synthetic and real image, using a discriminator. The output of the discriminator indicates the level of similarity between the real and generated content, which is subsequently used in a loss function to train both the generator and discriminator. With knowledge of the vectors that successfully generate realistic content similar vectors can be constructed to produce similar content. This introduces the idea of a latent space. Consider a 4-D input vector, $v=[0.2,0.1,0.99,0.27]$ with values between 0 and 1, that produces an image of a shoe. The values, also called latent variables, are used to infer various features. For instance, the first dimension, $v_{0}=0.2$ may be responsible for modelling features relating to shoe straps and laces, where a $v_{0}=0$ returns a laced shoe, $v_{0}=1$ returns a strapped shoe and $v_{0}=0.5$ returns a shoe with a strap and a lace. Pushing the analogy further, consider the relationship between different latent variables in our input vector. For example, if $v_{1}$ represents low-top or high-top shoes, then the combination of $v_{0}$ and $v_{1}$ will produce high-tops with laces that are tied at the tongue of the shoe and high-tops with laces that are tied at the throat of the shoe. This works so long as the training data set is sufficiently diverse to learn these various features. An acknowledged limitation for this type of GAN is the correlation between the dimensionality of the input vectors (which increases with the number of features/objects to be modelled) and computation cost (including training time, hardware and energy consumed).
To overcome this problem, conditional GANs [264, 176, 236, 121] condition the random input vectors on context-specific cues, such as object classes or scene descriptors. This is accomplished by inputting the additional information into the generator and discriminator, thus facilitates the generation of highly diverse content. For example, to generate images of wild animals, the additional information could be the class of animal (e.g. frog, bear and lion).
Stable Diffusion embodies a different approach to image and video content generation. Until recently, generative methods, such as GANs, encoder-decoders and autoencoders (described in Figure 2) have struggled to provide solutions that offer both low cost (energy, computation, etc.) and fast inference. This is the accomplishment of stable diffusion and a predominant reason for its current success. Stable Diffusion [7] works by first generating a random N-dimensional tensor - similar to the random vector generated in a GAN. The tensor, together with an encoded prompt, are input into a noise predictor (usually a U-Net [216, 231]) which outputs a new tensor predicting the noise in the input tensor. This noise tensor is subtracted from the input tensor, essentially denoising the input without reducing the resolution of the input tensor. The new tensor is re-used as an input (now without a prompt) and a new, less noisy, tensor is determined. This process repeats for a fixed number of steps, whereby the final and least noisy tensor is decoded into an image. As the step from a latent space to a less noisy space is a simple operation for a GPU to perform (i.e. a subtraction of two tensors), the generation process is fast. Additionally, with sufficient iterations, the generated image contains negligible noise and is thus of high quality.
In conclusion, generative AI is a powerful tool and may offer numerous solutions to production challenges moving forward. However, with great power researchers must actively seek ethical solutions to releasing code and/or pre-trained models. While many publicly available models (or online application interfaces) incorporate safety measures into their code (like detecting malicious text-based inputs) there are still cases where this is not sufficient for avoiding pornographic content generation [214]. For researchers, steps should be taken to raise and act on concerns prior to releasing publicly available code, dataset or pre-trainedd models.
@@ 2.3 Object Detection
In many instances in IC, there’s a need not just for automatic object recognition, but also for locating these objects within an image or a video sequence. For example, in live football broadcasting, there is a need to classify player poses for the purpose of automating camera control. This can assist upcoming phases of a game or event, leading to better organisation of shots. Considering the variation of background clutter (e.g., from a crowd, grass, dynamic advertisement banners, etc.), object detection poses significant challenges. Location can be identified by either pixel-level contouring around the object’s edge or by a bounding box. The former involves pixel-level classification and segmentation, together called semantic segmentation. In contrast, while the latter incorporates a regression branch to estimate the four corners of the bounding box alongside the classification branch. Object detection using bounding boxes is faster than semantic segmentation and has been extensively utilized in real-time applications, such as sports broadcasting. For example You Only Live Once (YOLO) models for object detection and image segmentation [207] only require one pass of a network to infer location. Furthermore, image masking can help YOLO reduce stationary noise [288] by excluding known pixels regions that do not contain relevant information during inference.
@@
YOLO works by dividing a target image/frame into $N$ cells having size $s\times s$ resolution; these are called residual blocks. Then, using bounding box regression the model estimates which residual blocks are associated with which bounding boxes. The probability of an object being present within the residual block is also predicted. Using non maximal suppression the model is able to reduce the influence of low probability scores and determine the bounding boxes with highest confidence; as illustrated in Figure 3. Following this, to measure the accuracy we take the intersection over union (IoU) of a ground truth bounding box and the estimated box to determine how close to the correct bounding box the prediction is. This provides a loss function to use during model optimisation Bandyopadhyay [15].
Figure 3: A visualisation of the stages of the YOLO process with an example of how the IoU is determined.
@@
Several versions of YOLO have been used in IC applications, the current variant being YOLOv8 [117]. YOLOv2 supports the prediction of a fixed number objects in a single block, resulting in improved detection of smaller and grouped objects. YOLOv3 builds on this by introducing a model (with higher complexity) that can detect smaller objects and preserve fine detail. YOLOv4 [22] (proposed by different authors) adds to YOLOv3 with cross mini batch normalisation for higher accuracy and weighted residual connections for better convergence during learning.
@@ 2.4 Camera Pose Estimation
Camera pose estimation approximates the orientation, path and motion (the extrinsics) of a camera in 3-D space which, for example, is useful as a basis for mitigating motion blur caused by the motion of a camera. Accurate camera tracking also enables automated object detection, human pose detection and 3-D capture (photogrammetry). There exist numerous (commercial) solutions for this involving additional hardware (e.g. LIDAR or GPS) for local and global positioning. Though, there is also research on video-based solutions for cases where physical positioning systems are unavailable.
There are three main approaches to pose estimation: (1) Visual Odometry (VO), (2) visual Simultaneous Localisation and Mapping (vSLAM) and (3) Structure from Motion (SfM), all of which share common components [241]. The objective of VO is to recover a camera path incrementally, optimizing the current pose given the prior set of poses (window bundle adjustment) [40]. We consider VO as a short-sighted solution as it has trouble linking a poses that intersect previously visited locations, i.e. the new set of frames would be treated as a separate locations. Thus, vSLAM learns an additional global localisation map as well as inheriting VO’s objective of optimising the consistency of the local trajectory. The constraint with using vSLAM is that there are limited number of cameras and window bundle adjustments that can be processed. SfM approaches [2, 58] overcome this by utilizing measurements taken for every viewpoint/frame to boost reliability and avoid degenerative cases [194], including those in Molton and Brady [168], Mallet et al [159] and Hirschmuller et al [100]. This is a popular means of pre-processing 2D images for 3-D modelling.

### v3

# Intelligent Cinematography: A review of AI research for cinematographic production

@@ Abstract
This paper offers the first comprehensive review of artificial intelligence (AI) research in the context of real camera content acquisition for entertainment purposes and is aimed at both researchers and cinematographers. Addressing the lack of review papers in the field of intelligent cinematography (IC) and the breadth of related computer vision research, we present a holistic view of the IC landscape while providing technical insight, important for experts across disciplines. We provide technical background on generative AI, object detection, automated camera calibration and 3-D content acquisition, with references to assist non-technical readers. The application sections categorize work in terms of four production types: General Production, Virtual Production, Live Production and Aerial Production. Within each application section, we (1) sub-classify work according to research topic and (2) describe the trends and challenges relevant to each type of production. In the final chapter, we address the greater scope of IC research and summarize the significant potential of this area to influence the creative industries sector. We suggest that work relating to virtual production has the greatest potential to impact other mediums of production, driven by the growing interest in LED volumes/stages for in-camera virtual effects (ICVFX) and automated 3-D capture for virtual modeling of real world scenes and actors. We also address ethical and legal concerns regarding the use of creative AI that impact on artists, actors, technologists and the general public.
@@
Creative Industries, Computer Vision, Machine Learning, Video Processing and Applications
@@ 1 Introduction
Intelligent cinematography (IC) leverages artificial intelligence (AI) to assist camera-based tasks throughout the media production workflow. For pre-production, this concerns shot planning, for example using a script ([224, 72]) or user input ([57, 285]). For production, this includes automated camera-control ([155, 62, 168, 26, 290]), methods for content acquisition ([176, 207, 309, 149]) and assisting directors and camera operators ([76, 156, 113, 111, 37, 28]). For post-production, this involves in-camera virtual effects (ICVFX) ([240, 134, 80, 187]), dealing with camera artefacts such as reflections ([283]), de-rigging and novel viewpoint synthesis ([135, 176, 296]).
Intelligent cinematography influences the worlds of cinema, gaming and televised broadcasting and streaming. The symbiotic relationship between research and video production has been historically important to the progression of video entertainment. Yet, as this relationship has evolved, the types of content and the styles of production have increasingly relied on a diverse pool of research across many domains. For example, we find work that automates manual capture processes ([72, 149]) as well as work that understands and enhances audience immersion ([113, 269]). This diversity and fluidity of activity makes it challenging to establish a unified definition and description of the IC research landscape. Hence there is an evident absence of, and need for, both general review papers and more specialized articles. In this paper, we attempt to overcome these challenges by defining and presenting a structured review on the state and future of IC research. Our definition of IC however specifically focuses on real content acquisition. We exclude research aimed at assisting post-production or automating content generation (e.g., using diffusion models) for two main reasons. Firstly, AI research relating to post-production is already well supported by review papers, like [85, 294, 241, 40]. Secondly, post-production AI typically deals with virtual cameras, objects and worlds, which are typically irrelevant for real content acquisition.
This paper categorizes work by production medium. Similar to the definition of an artistic medium, this groups research by the relevant tools and practices involved with the various production types, rather than grouping work by objective. For example, live sport and event broadcasting share many cinematographic practices, so are grouped under the same heading, Live Production. This allows us to: (1) further sub-categorize according to related AI research fields and provide a holistic view of the IC landscape while maintaining relevance to each production medium; and (2) include research topics which, while not targeted at creative video production, have direct relevance to it. The topics covered for each medium are outlined below and are visualized in Figure 1.
Figure 1: A taxonomy of IC research, including work within and outside the scope of this paper
@@
Scope and related work
Our focus is on camera-oriented research; hence we do not cover popular creative AI topics such as large language models (LLMs) for script generation or generative AI for video generation. We acknowledge, however, that these may directly impact cinematography in the future. For instance, LLMs may prove useful for text-driven camera control or workflow optimization. Moreover, the rapid progression of generative AI models could deliver high quality video-based methods that offer an alternative solution to automated 3-D capture. As the relevant literature does not yet consider camera-based objectives tied to these technologies we do not investigate these paradigms in the main body. Instead, we highlight potential links in Figure 1, provide background on generative AI in Appendix A.4 and discuss future use cases of these technologies within the relevant subsections.
There exist a small number of articles that review AI in the context of cinematography. For example, Christie et al [59] reviews camera control for computer graphics. This paradigm involves viewpoint computation and motion planning systems, and considers technical constraints such as direct and assisted control, and cinematic constraints (such as view composition and pose). Similarly, Chen and Carr [48] reviews autonomous methods for camera planning, control and selection for multi-camera systems. Furthermore, Qi et al [209] reviews camera calibration techniques. Other works look at use-cases for IC such as Galvane et al [85] for video editing, Young et al [294] and Sharma et al [241] for gaming and animation, and Mademlis et al [166] for unmanned aerial vehicle (UAV) control. Our survey aims to bridge related works to provide a holistic view of the IC landscape so that computing and cinematographic communities mutually benefit from the discussion.
It is important to note that this review was conducted between September 2022 and May 2024. Considering that many aspects of IC research continues to evolve at a rapid pace, the research represented in this paper will not always reflect the “current” state-of-the-art. Our intention is instead, to present an organized view of the IC landscape and envision potential areas for future research. In this way, researchers and industrialists alike may benefit, even in the longer term.
The structure of this paper is as follows. In Section 2 we introduce the technical background on AI research commonly found in IC. The main body is contained in Section 3, where the order of the subsections is reflected in the prior list. For each subsection, we present an overview of the current paradigms and approaches and conclude with a discussion on the current challenges and future works. Additionally, we append Section 3.5 to the main body to highlight the social, ethical and legal challenges associated with IC. Lastly, in Section 4, we offer concluding remarks and outline research that we believe has the most potential moving forward.
@@ 2.1 Convolutional Neural Networks
Convolutional neural networks (CNNs) are the primary technology behind IC. CNNs are a class of neural network that extract the features of an image-type by filtering through a neighborhood of pixels with learnable filter coefficients, to learn common patterns within a set of training images. In the context of IC, CNNs provide backbone architectures for numerous tasks, such as visual generative AI (Appendix A.4), object detection (Subsection 2.2) and aerial cinematography (Section 3.4).
A common application of CNNs is in image classification; the basic structure is shown in Figure 2. First, an image is input to the first layer through $N$ channels, where each channel represents a different representation of the image. For example, an RGB image will have three values (red, green and blue) for each pixel, thus $N=3$. The outputs of this layer then feed into a series of hidden layers. The outputs of each layer are called feature maps. The last hidden layer transforms these feature maps into feature vectors, so that the output layer can interpret a class probability using the feature vector. An image is then classified if its probability falls within a predefined confidence-interval. In the case of Figure 2, if our confidence interval is $0.9$, we classify the image as containing dogs.
@@
    A convolution layer applies an $n\times m$ kernel (filter) over each pixel neighborhood from an input channel. To deal with pixels on the perimeter of an image, we generally apply padding (the type of padding can vary). Dilated [297] or deformable convolutions [66] are also used to modify the structure of the kernel for special cases. In addition, a stride can be applied to skip some pixels, i.e. the filter applies to the pixels located $s$ distances from each-other ($s=1$ means all pixels are filtered and $s=2$ means every second pixel in a row and column is taken).
@@
    An activation function determines how the features of a layer are transformed into an output. To be used in a hidden layer, the activation function needs to be differentiable and nonlinear, otherwise we would not be able to calculate the gradients of parameters during a backward pass through the network. Rectified linear unit (ReLU) and LeakyReLUs layers are frequently used as they are a differentiable non-linear function, with the added benefit that they are not susceptible to vanishing gradients (where gradients of parameters become too small to make regressive change to the model’s parameters). Several types of activation function exist, as discussed in Baheti [11].
@@
    A fully-connected layer is one where each node is dependent on all the outputs from the previous layer. This allows the network to introduce a wide range of dependencies between parameters.
@@ 2.2 Object Detection
In many instances in IC, there is a need not just for automatic object recognition, but also for locating these objects within an image or a video sequence. For example, in live football broadcasting, there is a need to classify player poses for the purpose of automating camera control. This can assist upcoming phases of a game or event, leading to better organization of shots. Considering the variation of background clutter (e.g., from a crowd, grass, dynamic advertisement banners, etc.), object detection presents significant challenges. Location can be identified by either pixel-level contouring around the object’s edge or by a bounding box. The former involves pixel-level classification and segmentation, together called semantic segmentation. In contrast, while the latter incorporates a regression branch to estimate the four corners of the bounding box alongside the classification branch. Object detection using bounding boxes is faster than semantic segmentation and has been extensively utilized in real-time applications, such as sports broadcasting. For example You Only Live Once (YOLO) models for object detection and image segmentation [218] only require one pass of a network to infer location. Furthermore, image masking can help YOLO reduce stationary noise [301] by excluding known pixels regions that do not contain relevant information during inference. A real use-case for this was demonstrated by researchers at Kaunas University of Technology, who applied real-time online object detection and segmentation as a means of removing in-camera distractors for broadcasting a soccer game, KTU [140].
@@
YOLO works by dividing a target image/frame into $N$ cells having size $s\times s$ resolution; these are called residual blocks. Then, using bounding box regression the model estimates which residual blocks are associated with which bounding boxes. The probability of an object being present within the residual block is also predicted. Using non maximal suppression the model is able to reduce the influence of low probability scores and determine the bounding boxes with highest confidence; as illustrated in Figure 3. Following this, to measure the accuracy we take the intersection over union (IoU) of a ground truth bounding box and the estimated box to determine how close to the correct bounding box the prediction is. This provides a loss function to use during model optimization Bandyopadhyay [14].
Figure 3: Visualizing the YOLO process with an example of how the IoU is determined. The ground truth is represented by grey boxes and the predictions are the multi-colored bounding boxes
@@
Several versions of YOLO have been used in IC applications, the current variant being YOLOv8 [124], which was used in KTU [140] for detecting and segmenting distractors for soccer games. YOLOv2 supports the prediction of a fixed number objects in a single block, resulting in improved detection of smaller and grouped objects. YOLOv3 builds on this by introducing a model (with higher complexity) that can detect smaller objects and preserve fine detail. YOLOv4 [22] (proposed by different authors) adds to YOLOv3 with cross mini batch normalization for higher accuracy and weighted residual connections for better convergence during learning. Regarding IC, we find YOLO applied in live broadcasting (Section 3.3) and aerial cinematography (Section 3.4).
@@ 2.3 Camera Pose Estimation
Camera pose estimation approximates the orientation, path and motion (the extrinsics) of a camera in 3-D space which, for example, is useful as a basis for mitigating motion blur caused by the motion of a camera. Accurate camera tracking also enables automated object detection, human pose detection and 3-D capture (photogrammetry). There exist numerous (commercial) solutions for this involving additional hardware (e.g., LIDAR or GPS) for local and global positioning. Though, there is also research on video-based solutions for cases where physical positioning systems are unavailable.
There are three main approaches to pose estimation: (1) Visual Odometry (VO), (2) visual Simultaneous Localization and Mapping (vSLAM) and (3) Structure from Motion (SfM), all of which share common components [253]. The objective of VO is to recover a camera path incrementally, optimizing the current pose given the prior set of poses (window bundle adjustment) [42]. We consider VO as a short-sighted solution as it has trouble linking a poses that intersect previously visited locations, i.e. the new set of frames would be treated as a separate locations. Thus, vSLAM learns an additional global localization map as well as inheriting VO’s objective of optimizing the consistency of the local trajectory. The constraint with using vSLAM is that there are limited number of cameras and window bundle adjustments that can be processed. SfM approaches [2, 60] overcome this by utilizing measurements taken for every viewpoint/frame to boost reliability and avoid degenerative cases [204], including those in Molton and Brady [178], Mallet et al [169] and Hirschmuller et al [107]. This is a popular means of pre-processing 2D images for 3-D modeling.